# NAAQA: A Neural Architecture for Acoustic Question Answering

### v1

@@ Abstract
The goal of the Acoustic Question Answering (AQA) task is to answer a free-form text question about the content of an acoustic scene. It was inspired by the Visual Question Answering (VQA) task. In this paper, based on the previously introduced CLEAR dataset, we propose a new benchmark for AQA, that emphasizes the specific challenges of acoustic inputs, e.g. variable duration scenes. We also introduce NAAQA, a neural architecture that leverages specific properties of acoustic inputs. The usage of time and frequency 1D convolutions to process 2D spectro-temporal representations of acoustic content shows promising results and enables reductions in model complexity. NAAQA achieves 91.6% of accuracy on the AQA task with $\sim$7 times fewer parameters than the previously explored VQA model. We provide a detailed analysis of the results for the different question types. The effectiveness of coordinate maps in this acoustic context was also studied and we show that time coordinate maps augment temporal localization capabilities which enhance performance of the network by $\sim$17 percentage points.
@@ 1 Introduction
Question answering (QA) tasks are examples of constrained and limited scenarios for research in reasoning. The agent’s task in QA is to answer questions based on context. Text-based QA uses text corpora as context [undef, undefa, undefb, undefc, undefd, undefe]. In visual question answering (VQA) the questions are related to a scene depicted in still images [undeff, undefg, undefh, undefi, undefj, undefk, undefl]. Finally, video question answering attempts to use both the visual and acoustic information in video material as context [undefm, undefn, undefo, undefp, undefq, undefr]. However, the use of the acoustic channel is usually limited to linguistic information that is expressed in text form, either with manual transcriptions (e.g. subtitles) or by automatic speech recognition [undefs].
@@
The audio modality provides important information that has not yet been exploited in QA reasoning. Audio allows QA systems to answer relevant questions more accurately, or even to answer questions that are not approachable from the visual domain alone. In [undefy] we introduced, to our knowledge, the first task in acoustic question answering. The agent’s goal was to answer questions related to acoustic scenes composed by a sequence of elementary sounds. The questions foster reasoning on the properties of the elementary sounds and their relative and absolute position in the scene. As a benchmark for this task, we created the CLEAR dataset.
The modeling in [undefy] was mutated from a similar task in VQA. However, the structure of acoustic data is fundamentally different from that of visual data. This is illustrated in [undefz] where two standard data sets in computer vision (MNIST) and speech technology (Google Speech Commands) are compared via T-SNE [undefaa].
In this paper we address the specific nature of acoustic observations in the context of Acoustic Question Answering (AQA) by introducing a number of extensions and improvements both to the data and to the methods to solve the AQA task. The main contributions are summarized as follows: we propose
  • a neural architecture that leverages specific properties of acoustic inputs,
  • a more challenging version of the CLEAR dataset which comprises scenes of variable duration (compared to fixed duration in [undefy]),
  • an analysis of the relevance of coordinate maps in an acoustic context, and
  • a detailed comparison of the performance of the Visual FiLM [undefab] network initially designed for the VQA task with the substantially smaller proposed NAAQA model.
The rest of the paper is organized as follows: Section 2 reports on recent related work, Section 3 describes our dataset, Section 4 presents the QA models we have tested, Section 5 gives details on the experimental settings, Section 6 discusses the results and, finally, Section 7 concludes the paper.
@@ 2.2 Visual Question Answering (VQA)
Visual Question Answering aims to answer questions based on a visual scene. Several VQA datasets are available to the scientific community [undeff, undefac, undefad, undefae, undefaf, undefag, undefh, undefi, undefah, undefg]. However, designing an unbiased dataset is non-trivial. [undefj] [undefj] observed that the type of questions has a strong impact on the results of neural network based systems which motivated research to reduce the bias in VQA datasets [undefai, undefaj, undefak, undefk, undefl, undeff]. Gathering good labeled data is also non-trivial which induced [undefk] and [undefl] [undefk, undefl] to constrain their work to yes/no questions. To alleviate this problem, [undeff] [undeff] proposed the use of synthetic data for both questions and visual scenes. The resulting CLEVR dataset has been extensively used to evaluate neural networks for VQA applications [undefab, undefal, undefam, undefan, undefao, undefap] which helped foster research on VQA. To create visual scenes, the authors automated a 3D modelling software. This allows for an unlimited supply of labeled data eliminating the time and effort needed for manual annotations. For the questions, they first manually designed semantic representations for each type of question. These representations describe the reasoning steps needed to answer a question (i.e. ”find all cubes $|$ that are red $|$ and metallic”). The semantic representations are then instantiated based on the visual scene composition thus creating a question and an answer for a given scene. This gives complete control over the labelling process.
2.3 Acoustic Question Answering (AQA)
To the best of our knowledge, we were the first to introduce the AQA task in [undefy]. [undefaq] [undefaq] also presented reflections on acoustic reasoning via question answering. They proposed a network comprising 2 sets of FiLM layers [undefab] per Resblock to modulate the activation maps based on both textual and audio features contexts. This second set of FiLM layers makes this model much more complex in terms of number of parameters than the initial network proposed by [undefab] [undefab]. As one of our goal is to solve the CLEAR task with a minimal number of parameters, this model is out of scope for this study.
Using synthetic data in the design of a dataset has substantial advantages. Data can be automatically annotated which save time and complexity. The number of training examples that can be generated is only limited by the available computational resources. Controlling the generation process gives a complete understanding of the properties and relations of the objects in a scene. This understanding can be leveraged to reduce bias in the dataset and to generate complex questions and their corresponding answers. The CLEAR dataset is generated using semi-synthetic data with a methodology similar to [undeff]
@@ 2.4 Convolutional neural network on Audio
Convolutional neural networks (CNN) have dominated the visual domain in recent years. More recently, they have also been applied to a number of problems in the acoustic domains such as acoustic scene classification [undefar, undefas, undefat, undefau], music genre classification [undefav, undefat], instrument classification [undefaw, undefax], sound event classification and localization [undefay] and speech recognition [undefat]. Some authors [undefav, undefaw, undefax, undefau] use intermediate representations such as STFT [undefaz], MFCC [undefaaa] or CQT [undefaab] spectrograms while others [undefat, undefar] work directly with the raw audio signal.
@@
Square convolutional and pooling kernels are often used to solve visual task such as VQA, visual scene classification and object recognition[undefaac, undefaad, undefaae, undefaaf]. [undefas, undefaag, undefaah] [undefas, undefaag, undefaah] have successfully used visually motivated CNN with square filters to solve audio related tasks. Time-frequency representations of audio signals are however structured very differently than visual representations. [undefaai] [undefaai] explores the performance of different structures of convolutive kernels when working with audio signals. They propose the use of 1D convolution kernels to capture time-specific or frequency-specific features. They demonstrate that similar accuracy can be reached using a combination of 1D convolutions instead of 2D convolutions by combining 1D time and 1D frequency filters while using much fewer parameters. They also explore rectangular kernels which capture both time and frequency features at different scales.
In this work, a more challenging version of the CLEAR dataset which comprises more elementary sounds and scenes of variable duration (compared to fixed duration in [undefy]) is introduced. The performance of a network initially designed for the VQA task (Visual FiLM) [undefab] is evaluated on the AQA task, and modification to the architecture are proposed to leverage specific properties of acoustic inputs. The effectiveness of using coordinate maps in an acoustic context is also analyzed. Finally, the architecture is optimized to reduce complexity in terms of number of parameters.
3 Variable scene duration dataset
@@
In this section, we briefly review the principal aspects of the dataset and present an updated version for a task that emphasizes the challenges with AQA as opposed to VQA. A graphical overview of the CLEAR dataset generation process is depicted in Figure 1. Each record in the dataset is a unique combination of a scene, a question and an answer.
@@
To build acoustic scenes, we prepared a bank of elementary sounds composed of real musical instrument recordings extracted from the Good-Sounds [undefaaj] dataset. The bank comprises 135 unique recordings (compared to 56 in previous work [undefy]) sampled at 48 KHz including 6 different instruments (bass, cello, clarinet, flute, trumpet and violin), 12 notes (chromatic scale) and 3 octaves. The acoustic scenes are built by concatenating between 5 to 15 randomly chosen sounds from the elementary sound bank into a sequence (Previous work [undefy] comprised fixed duration scenes). Silence segments of random duration are added in-between elementary sounds. The acoustic scenes are then corrupted by filtering the scenes to simulate room reverberation and by adding a white uncorrelated uniform noise. Both additive noise and reverberation vary from scenes to scenes which reduces overfitting and provides a more natural acoustic scene.
Each elementary sound in a scene is characterized by an n-tuple: [Instrument, Brightness, Loudness, Musical Note, Duration, Absolute position in scene, Relative position in scene, Global position]. The Brightness property is computed by using the timbralmodels [undefaak] library. A threshold is used to define the label of the sound (Dark or Bright). The Loudness labels are assigned based on the perceptual loudness as defined by the ITU-R BS.1770-4 international normalization standard [undefaal]. Again, a threshold is used to determine if the sound is Quiet or Loud. All attribute values are listed in Table I as possible answers to the questions explained below.
@@
For each scene, a number of questions is generated using CLEVR-like [undeff] templates. A template defines the reasoning steps required to answer a question based on the composition of the scene (i.e. ”find all instances of violin $|$ that plays before trumpet $|$ that is the loudest”). 942 templates where designed for this AQA task. Not all template instantiation results in a valid question. The generated questions are filtered to remove ill posed questions similarly to [undeff]. Table I shows examples of questions with their answers.
@@
The a priori probability of answering correctly with no information about the question or the scene, and assuming a uniform distribution of classes, is $\frac{1}{57}=1.75\%$. These probabilities are lower, on average, if we introduce information about the question. For example, if we know that the question is of the type Exist or Counting comparison, there are only two possible answers (yes or no) and the probability of answering correctly by chance is $0.5$. The majority class accuracy (always answering the most common answer) is 7.5%. Some statistics on the CLEAR dataset are presented in Table II.
@@
The generation process was built with extensibility in mind. Different versions of the dataset with fewer or more objects per scene can be generated by using different parameters for the generation script. It is also possible to modify the elementary sound bank to generate datasets for AQA in other domains, speech or environmental sounds, for example. The code for generating the dataset is available on GitHub [^1] The link will be made available upon publication.
Dataset global statistics
Nb of questions 200 000 Nb of scenes 50 000 Nb of unique vocabulary words 91
Dataset detailed statistics
@@
Mean Min Max Nb of sounds per scene 10 5 15 Elementary sound duration 0.85s 0.69s 1.11s Scene duration 10.69s 4.46s 17.82s Nb of words per question 17 6 28 Nb of unique words per question 12 5 19
TABLE II: Dataset statistics

### v2

@@ Abstract
The goal of the Acoustic Question Answering (AQA) task is to answer a free-form text question about the content of an acoustic scene. It was inspired by the Visual Question Answering (VQA) task. In this paper, based on the previously introduced CLEAR dataset, we propose a new benchmark for AQA, namely CLEAR2, that emphasizes the specific challenges of acoustic inputs. These include handling of variable duration scenes, and scenes built with elementary sounds that differ between training and test set. We also introduce NAAQA, a neural architecture that leverages specific properties of acoustic inputs. The use of 1D convolutions in time and frequency to process 2D spectro-temporal representations of acoustic content shows promising results and enables reductions in model complexity. We show that time coordinate maps augment temporal localization capabilities which enhance performance of the network by $\sim$17 percentage points. On the other hand, frequency coordinate maps have little influence on this task. NAAQA achieves 79.5% of accuracy on the AQA task with $\sim$4 times fewer parameters than the previously explored VQA model. We evaluate the perfomance of NAAQA on an independent data set reconstructed from DAQA. We also test the addition of a MALiMo module in our model on both CLEAR2 and DAQA. We provide a detailed analysis of the results for the different question types. We release the code to produce CLEAR2 as well as NAAQA to foster research in this newly emerging machine learning task.
@@ 1 Introduction
Question answering (QA) tasks are examples of constrained and limited scenarios for research in reasoning. The agent’s task in QA is to answer questions based on context. Text-based QA uses text corpora as context [undef, undefa, undefb, undefc, undefd, undefe]. In visual question answering (VQA) the questions are related to a scene depicted in still images [undeff, undefg, undefh, undefi, undefj, undefk, undefl]. Finally, video question answering attempts to use both the visual and acoustic information in video material as context [undefm, undefn, undefo, undefp, undefq, undefr]. The use of the acoustic channel is usually limited to linguistic information that is expressed in text form, either with manual transcriptions (e.g. subtitles) or by automatic speech recognition [undefs].
@@
The audio modality provides important information that can be leveraged in the context of QA reasoning. Audio allows QA systems to answer relevant questions more accurately, or even to answer questions that are not approachable from the visual domain alone. In [undefy], we introduced the AQA task and proposed a new database (CLEAR) to promote research in AQA. The agent’s goal, in the proposed task, was to answer questions related to acoustic scenes composed by a sequence of elementary musical sounds. The questions foster reasoning on the properties of the elementary sounds and their relative and absolute position in the scene. To build CLEAR, we were inspired by the work of Johnson et al. [undeff] for VQA. Similarly, we tested an architecture built for VQA and based on FiLM layers [undefz] on the newly proposed AQA task. Next, [undefaa] [undefaa] proposed to extend the questions to more acoustically realistic situations by developing a new database called DAQA. To evaluate the results, they proposed the MALiMo network which relies on several FiLM layers.
The works cited above use neural network architectures that are largely inspired by image processing research. However, the structure of acoustic data is fundamentally different from that of visual data. This is illustrated for example in [undefab] where two standard data sets in computer vision (MNIST) and speech technology (Google Speech Commands) are compared via T-SNE [undefac]. A legitimate question is whether it is possible to obtain better results (in terms of accuracy and network complexity) by adapting the first layers of the architectures to take into account intrinsic characteristics of acoustic signals. Even within the AQA domain, the properties of acoustic data may vary significantly depending on the nature of the auditory scenes (e.g. CLEAR vs DAQA). It is, therefore interesting to evaluate the impact of the dataset on system performance.
To answer the above questions, we present a study that evaluates the impact of audio pre-processing, of acoustic feature extraction and of dataset characteristics on the performance neural architectures for AQA. When considering performance, we focus both on accuracy and complexity of the models. The main contributions can be summarized as follows:
  • We introduce CLEAR2 a more challenging version of the CLEAR dataset, which comprises scenes of variable duration and different elementary sounds for the training and test sets,
  • we regenerate a version of the DAQA dataset (DAQA)
  • we evaluate different ways of using a VQA architecture on AQA inputs,
  • we study the effect of time and frequency coordinate maps for acoustic data at different levels in the architecture,
  • we introduce new feature extraction modules motivated by the properties of acoustic data,
  • based on the above points, we propose a highly optimized FiLM-based architecture (NAAQA) that outperforms models that are 5 to 10 times more complex in terms of number of parameters,
  • we test the addition of a MALiMo module (from [undefaa]) into our NAAQA architecture,
  • we compare the performance NAAQA and NAAQA+MALiMo on CLEAR2 and DAQA’ to test generality of the methods,
  • we provide a detailed analysis of our results based on question type to improve interpretability.
The rest of the paper is organized as follows: Section 2 reports on recent related work, Section 3 describes both our CLEAR2 dataset and the DAQA dataset, Section 4 presents the QA models we have tested, Section 5 gives details on the experimental settings, Section 6 presents and discusses the results and, finally, Section 7 concludes the paper. Some extra information is reported in the supplementary material, and referred to in the paper through the corresponding section.
@@ 2 Related Work
This section presents previous research in QA systems and the needs for databases in AQA research.
@@ 2.2 Visual Question Answering (VQA)
Visual Question Answering aims to answer questions based on a visual scene. Several VQA datasets are available to the scientific community [undeff, undefad, undefae, undefaf, undefag, undefah, undefh, undefi, undefai, undefg]. However, designing an unbiased dataset is non-trivial. [undefj] [undefj] observed that the type of questions has a strong impact on the results of neural network based systems which motivated research to reduce the bias in VQA datasets [undefaj, undefak, undefal, undefk, undefl, undeff]. Gathering good labeled data is also non-trivial which induced [undefk] and [undefl] [undefk, undefl] to constrain their work to yes/no questions. To alleviate this problem, [undeff] [undeff] proposed the use of synthetic data for both questions and visual scenes. The resulting CLEVR dataset has been extensively used to evaluate neural networks for VQA applications [undefz, undefam, undefan, undefao, undefap, undefaq] which helped foster research on VQA. To create visual scenes, the authors automated a 3D modelling software. This allows for an unlimited supply of labeled data eliminating the time and effort needed for manual annotations. For the questions, they first manually designed semantic representations for each type of question. These representations describe the reasoning steps needed to answer a question (i.e. “find all cubes $|$ that are red $|$ and metallic”). The semantic representations are then instantiated based on the visual scene composition thus creating a question and an answer for a given scene. This gives complete control over the labelling process.
2.3 Dababases for AQA
As in VQA, using generated data in the design of AQA datasets has substantial advantages. Data can be automatically annotated which saves time and complexity. The number of training examples that can be generated is only limited by the available computational resources. Controlling the generation process gives a complete understanding of the properties and relations of the objects in a scene. This understanding can be leveraged to reduce bias in the dataset and to generate complex questions and their corresponding answers. The CLEAR dataset  [undefy] has been initially generated using semi-synthetic data. The elementary sounds were recordings of musical notes played by various instruments and players. However, the auditory scenes obtained by combining the elementary sounds had fixed duration. Moreover, the same elementary sounds were used to generate the test and training scenes. The DAQA dataset [undefaa] comprises more complex and less stationary elementary natural sounds coming for example from aircrafts, cars, doors, human speaking, bird singing, dog barking, etc. Although more complex and varied than CLEAR, the evaluation also uses the same elementary sound recordings for training and testing.
In this paper, we propose a more challenging version of CLEAR which uses different elementary sound recordings for the training and test sets and generates variable duration auditory scenes.
@@ 2.4 Convolutional neural network on Audio
Convolutional neural networks (CNN) have dominated the visual domain in recent years. More recently, they have also been applied to a number of problems in the acoustic domains such as acoustic scene classification [undefar, undefas, undefat, undefau], music genre classification [undefav, undefat], instrument classification [undefaw, undefax], sound event classification and localization [undefay] and speech recognition [undefat]. Some authors [undefav, undefaw, undefax, undefau] use intermediate representations such as STFT [undefaz], MFCC [undefaaa] or CQT [undefaab] spectrograms while others work directly with the raw audio signal [undefat, undefar].
@@
Square convolutional and pooling kernels are often used to solve visual task such as VQA, visual scene classification and object recognition [undefaac, undefaad, undefaae, undefaaf]. [undefas, undefaag, undefaah] [undefas, undefaag, undefaah] have successfully used visually motivated CNN with square filters to solve audio related tasks. Time-frequency representations of audio signals are however structured very differently than visual representations. [undefaai] [undefaai] explore the performance of different structures of convolutive kernels when working with music signals classification. They propose the use of 1D convolution kernels to capture time-specific or frequency-specific features. They demonstrate that similar accuracy can be reached using a combination of 1D convolutions instead of 2D convolutions by combining 1D time and 1D frequency filters while using much fewer parameters. They also explore rectangular kernels which capture both time and frequency features at different scales. The impact of such strategies for music classification is still an open question in the context of auditory scene analysis.
Coordinate maps initially proposed in [undefaaj] by [undefaaj] have proven successful for processing visual data in the context of VQA. The method consists in augmenting the visual input with matrices containing numbers in the range -1 to 1 which vary either in the $x$ or in the $y$-dimension. With MALiMo [undefaa], the same strategy is used to indicate the simultaneous relative positions of features in frequency and time. [undefaak] [undefaak] proposed Frequency-Aware convolutions which are equivalent to concatenating coordinate maps only in the frequency axis. The effectiveness of coordinate maps on the time dimension for audio signals have not been studied to the best of our knowledge.
In this study we first evaluate the performance of a network initially designed for the VQA task (Visual FiLM) [undefz] on the AQA task. Then we introduce the NAAQA architecture to leverage specific properties of acoustic inputs. For this architecture, we analyze the influence of separate time and frequency coordinate maps. We then study the impact of adding a MALiMo block into our architecture. Finally, we evaluate our model on the DAQA dataset.
3 Data
We use two very different datasets in our experiments in order to study the effect of the AQA task characteristics on the model performance. The first set, that is also a contribution of this paper, comprises musical sounds (CLEAR2); the second includes short environmental sounds (DAQA).
3.1 CLEAR2
@@
CLEAR2 is an updated version of CLEAR [undefy]. A graphical overview of the generation process is depicted in Figure 1. Each record in the dataset is a unique combination of a scene, a question and an answer.
@@
To build acoustic scenes, we prepared a bank of elementary sounds composed of real musical instrument recordings extracted from the Good-Sounds [undefaal] dataset [^1] Each elementary sound in a scene is characterized by an n-tuple: [Instrument, Brightness, Loudness, Musical Note, Duration, Absolute position in scene, Relative position in scene, Global position]. The Brightness property is computed by using the timbralmodels [undefaam] library. A threshold is used to define the label of the sound (Dark or Bright). The Loudness labels are assigned based on the perceptual loudness as defined by the ITU-R BS.1770-4 international normalization standard [undefaan]. Again, a threshold is used to determine if the sound is Quiet or Loud. All attribute values are listed in Table I as possible answers to the questions explained below.. Differently from CLEAR, in CLEAR2 we make sure that the recordings (players, instruments, microphones) of the elementary sounds are different for the training and test sets. For the training set, the bank comprises 135 unique recordings (compared to 56 in CLEAR) sampled at 48KHz including 6 different instruments (bass, cello, clarinet, flute, trumpet and violin), 12 notes (chromatic scale) and 3 octaves. A different set of 135 recordings of the same instruments recorded using different microphones and players is used to create the test set. The acoustic scenes are built by concatenating between 5 to 15 randomly chosen sounds from the elementary sound bank into a sequence (as opposed to CLEAR which comprised fixed duration scenes). Silence segments of random duration are added in-between elementary sounds. The acoustic scenes are then corrupted by filtering to simulate room reverberation and by adding a white uncorrelated uniform noise. Both the amount of noise and reverberation vary from scene to scene with the goal of reducing overfitting.
@@
For each scene, a number of questions is generated using CLEVR-like [undeff] templates. A template defines the reasoning steps required to answer a question based on the composition of the scene (i.e. “find all instances of violin $|$ that plays before trumpet $|$ that is the loudest”). 942 templates where designed for this AQA task. Not all template instantiations results in a valid question. The generated questions are filtered to remove ill posed questions similarly to [undeff]. Table I shows examples of questions with their answers.
@@
The a priori probability of answering correctly with no information about the question or the scene, and assuming a uniform distribution of classes, is $\frac{1}{57}=1.75\%$. These probabilities are higher, on average, if we introduce information about the question. For example, if we know that the question is of the type Exist or Counting comparison, there are only two possible answers (yes or no) and the probability of answering correctly by chance is $0.5$. The majority class accuracy (always answering the most common answer: Yes) is 7.5%. Statistics on the CLEAR2 dataset are presented in Table II.
@@
The generation process was built with extensibility in mind. Different versions of the dataset with fewer or more objects per scene can be generated by using different parameters for the generation script. It is also possible to modify the elementary sounds bank to generate datasets for AQA in other domains, speech or environmental sounds, for example. The code for generating the dataset is available on GitHub [^2] The link will be made available upon publication.
Datasets global statistics
Dataset CLEAR2 DAQA # of questions 200 000 599 441 # of scenes 50 000 100 000 # of answers 57 52 # of elementary sounds 135 ( + 135 for test) 358 # of types of question 11 5 # of unique vocabulary words 91 158
CLEAR2 Dataset detailed statistics
@@
Mean Min Max # of sounds per scene 10 5 15 Elementary sound duration 0.85s 0.69s 1.11s Scene duration 10.69s 4.46s 17.82s # of words per question 17 6 28 # of unique words per question 12 5 19
DAQA’ Dataset detailed statistics
Mean Min Max # of sounds per scene 9 5 12 Elementary sound duration 9.35s 0.6s 20s Scene duration 1min 19s 9s 3min 4s # of words per question 13 5 27 # of unique words per question 11 5 22
TABLE II: Datasets statistics
3.2 Reproducing DAQA
We were not able to fully recreate the DAQA dataset because it relies on some AudioSet [undefaao] youtube videos that have since been deleted. We were able to retrieve 358 sounds out of the 400 sounds that were used to generate the original dataset. We used these sounds to generate the dataset. Changing the number of elementary sounds also impacts the whole generation process. This dataset is therefore different from the original DAQA and will be referred to as DAQA from now on. Our results are therefore not fully comparable to the ones reported in [undefaa]. A list of all the missing sounds is available in Supplementary G.
3.3 Comparing CLEAR2 and DAQA
Table II report statistics on both CLEAR2 and DAQA. The major difference between both datasets is the type of elementary sounds used to generate the acoustic scenes, that is, sustained musical notes (CLEAR2) versus possibily transient environmental sounds (DAQA).
Acoustic scenes in DAQA are much longer on average than the ones in CLEAR2. This results in much bigger input spectrograms, and, in turns, much higher computational requirements and longer training time. Besides computational costs, the reason for producing shorter acoustic scenes in CLEAR2 was to produce a task that could be solved by a human. It would be unreasonable to ask, e.g., if the “flute plays before the sixth violin” in a 2 minute audio clip.
Finally, the original DAQA [undefaa] and, consequently, our reconstruction (DAQA) suffer from the same problem as the original CLEAR. The same elementary sounds are used in the training and test scenes. Although scenes are still different between training and test set, this may cause the models to “remember” the elementary sounds rather than extracting their properties. In CLEAR2, this problem was mitigated by using different elementary sounds for the training and test set.